If 16x – 4 and 5^x+y = 625, then y equals ____
Found 2 solutions by ikleyn, Alan3354:
Answer by ikleyn(52878)

.

1.  The fragment  "16x – 4" is NEITHER equation NOR inequality, so it does not define any condition at all.



2.  The fragment "5^x+y = 625" is ambiguous, since it admits at least two different readings:


    a)  (5^x)+y = 625  or

    b)  5^(x+y) = 625.



    To make the formula unambiguous use parentheses.


    Remember:  when you post formulas to this forum,  primitive "copy and paste"  DOES NOT WORK.


When you try to get help from this forum and want to interact with the tutors effectively, it requires at least two things from you:

      Be accurate and attentive in preparing your post, and

      Understand what you are doing.


--------------

Also,  do not post your problem to me personally through the "Tnank you" window  (as you recently did),
since in such case I don't know to whom to answer.


The "Thank you" window is designed for different purposes.



Answer by Alan3354(69443)

Makes no sense.